### v1

# Some new cases of Zilber–Pink in $Y(1)^{3}$

Abstract.
We prove the Zilber–Pink conjecture for curves in $Y(1)^{3}$ that intersect a modular curve in the boundary. We also give an unconditional result for points having few places of supersingular reduction. Both results are proved using the G-function method for unlikely intersections.
1. Introduction
The Zilber–Pink conjecture is one of the central problems in arithmetic geometry today. It can be formulated in many different settings but, for (pure) Shimura varieties, we can state it as follows.
Conjecture 1 (Zilber–Pink).
Let $V$ be an irreducible subvariety of a Shimura variety $S$. Suppose that the intersection of $V$ with the union of all special subvarieties of $S$ of codimension greater than $\dim(V)$ is Zariski dense in $V$. Then $V$ is contained in a proper special subvariety of $S$.
The simplest Shimura variety is $Y(1)\cong\mathbb{A}^{1}$ —the moduli space of elliptic curves. Clearly, Zilber–Pink says nothing for $Y(1)$ itself. For $Y(1)^{2}$ the conjecture asserts that an irreducible (plane) curve containing infinitely many special (CM) points is special. This was proved by André [And98] in 1998, and is prototypical of the general André–Oort conjecture, which states that if an irreducible subvariety of a Shimura variety contains a Zariski dense set of special points, then it is a special subvariety. The André–Oort conjecture has recently been proved in full, through the works of many authors, culminating in a paper of Pila–Shankar–Tsimerman, with an appendix by Esnault–Groechenig [PST].
Thus, the obvious next step is Zilber–Pink in $Y(1)^{3}$. With André–Oort decided, this reduces to the statement that an irreducible curve $C$ in $\mathbb{A}^{3}$ not contained in any proper special subvariety intersects only finitely many special curves.
If $C$ is not defined over $\overline{\mathbb{Q}}$, then Zilber–Pink holds for $C$ by work of Pila [Pil17, Thm. 1.4]. Over $\overline{\mathbb{Q}}$, the work of Habegger–Pila [HP12, Thm. 2] establishes that $C$ will only intersect finitely many special curves with a fixed coordinate. We recall that the special subvarieties of $Y(1)^{n}$ are defined by taking irreducible components of those subvarieties defined by finite conjunctions of
  • (1)
    $X_{i}=x$, for some $x$ in the set $\Sigma_{\rm CM}$ of singular moduli;
  • (2)
    $\Phi_{N}(X_{i},X_{j})=0$, for some $1\leq i<j\leq n$ and $N\in\mathbb{N}$,
where $\Phi_{N}(X,Y)\in\mathbb{Z}[X,Y]$ denotes the $N$th modular polynomial, defined by the property that two elliptic curves with $j$-invariants $j_{1}$ and $j_{2}$ are isogenous, with minimal isogeny of degree $N$, if and only if $\Phi_{N}(j_{1},j_{2})=0$. Special subvarieties defined by conditions only of type (2)—namely, those with no fixed coordinate—are examples of so-called strongly special subvarieties [CU05].
It follows, then, that the problem of Zilber–Pink in $Y(1)^{3}$ is reduced to the following, rather elegant conjecture.
Conjecture 2.
Let $C$ be an irreducible algebraic curve in $Y(1)^{3}\cong\mathbb{A}^{3}$ defined over $\overline{\mathbb{Q}}$ not contained in a proper special subvariety. Then $C$ contains only finitely many points $(x_{1},x_{2},x_{3})$ for which there exist $M,N\in\mathbb{N}$ such that
$\Phi_{M}(x_{1},x_{2})=\Phi_{N}(x_{2},x_{3})=0.$
Note that, by the preceding discussion, it suffices to consider only those points $s=(x_{1},x_{2},x_{3})\in Y(1)^{3}$ whose coordinates are non-singular moduli. For such points, the $M,N\in\mathbb{N}$ for which
$\Phi_{M}(x_{1},x_{2})=\Phi_{N}(x_{2},x_{3})=0$
are uniquely determined. We henceforth refer to such a point as modular and define
$\Delta(s):=\max\{M,N\}.$
By an ingenious argument, Habegger–Pila show that Conjecture 2 holds when $C$ is asymmetric, which (in this case) is to say that the degrees of the coordinate projections $C\to Y(1)$ are not all equal [HP12, Thm. 1]. In some sense, this covers most curves. On the other hand, it is not clear to us how to generalize their argument, or how to deal with symmetric curves specifically.
With this in mind, the present authors have more recently explored an alternative approach, based on techniques of André and Bombieri using G-functions. In [DO22], the first and second authors showed that Conjecture 2 holds for any $C$ whose Zariski closure $\overline{C}$ in $X(1)^{3}\cong(\mathbb{P}^{1})^{3}$ (the Baily–Borel compactification) contains $(\infty,\infty,\infty)$. In [Pap], the third author observed that this could be extended to any curve $C$ for which $\overline{C}\setminus C$ contains a point $(x_{1},x_{2},x_{3})$ with $x_{1},x_{2},x_{3}\in\{\infty\}\cup\Sigma_{\rm CM}$. In other words, $\overline{C}$ intersects a special point in the boundary.
Note that the boundary can naturally be identified with the disjoint union
$Y(1)^{2}\sqcup Y(1)^{2}\sqcup Y(1)^{2}\sqcup Y(1)\sqcup Y(1)\sqcup Y(1)\sqcup\{\infty\}$
of products of modular curves and the point $\infty:=(\infty,\infty,\infty)$. Since $\overline{C}$ must always intersect the boundary, it is interesting to ask which intersection conditions facilitate our method. A natural next step in this direction is the case of a curve intersecting a special curve in the boundary.
To make this precise, we say that a point in the boundary is modular if it belongs to one of the copies of $Y(1)^{2}$ and its coordinates $x_{1},x_{2}$ satisfy $\Phi_{N}(x_{1},x_{2})=0$ for some $N\in\mathbb{N}$. In this terminology, we prove the following.
Theorem 3.
Conjecture 2 holds under the assumption that $\overline{C}\setminus C$ contains a modular point.
However, our methods also yield a Zilber–Pink–type result without conditions at the boundary. In order to state it, let $C$ be as in the statement of Conjecture 2. Observe that we can always assume that $C$ contains a modular point $(b_{1},b_{2},b_{3})$. (Otherwise, there is nothing to prove!) Observe also that any modular point on $C$ belongs to $C(\overline{\mathbb{Q}})$. Thus, we denote by $K_{0}$ the number field $\mathbb{Q}(b_{1},b_{2},b_{3})$, and we say that a finite place $v$ of any finite extension of $K_{0}$ is a place of supersingular reduction if the elliptic curve with $j$-invariant $b_{1}$ (equivalently, $b_{2}$ or $b_{3}$) has supersingular reduction at $v$ (for the minimal Weierstrass model over $K_{v}$). For any such place $v$, lying over rational prime $p$, we denote by $|\cdot|_{v}$ the absolute value extending the standard $p$-adic absolute value on $\mathbb{Q}$. For any $\delta>0$, we say that a modular point $s=(x_{1},x_{2},x_{3})\in C$ has supersingular exponent $\delta$ if the set of places $v$ of $K_{0}(x_{1},x_{2},x_{3})$ of supersingular reduction for which
$|x_{1}-b_{1}|_{v},|x_{2}-b_{2}|_{v},|x_{3}-b_{3}|_{v}<1$
has cardinality at most $\Delta(s)^{\delta}$.
Theorem 4.
Let $C$ and $(b_{1},b_{2},b_{3})$ be as above. There exists $\delta>0$ such that $C$ contains only finitely many modular points of supersingular exponent $\delta$.
To prove Theorems 3 and 4, we obtain the following heights bounds. The derivations of Theorems 3 and 4 from these height bounds is now standard—facilitated by the so-called Pila–Zannier strategy and Masser–Wüstholz-type isogeny estimates—but we will give the details briefly in Section 5. In order to state them, we denote by $h(x_{1},x_{2},x_{3})$ the maximum of the absolute logarithmic heights $h(x_{i})$, as defined, for example, in [HP12, §2.2].
Theorem 5.
Let $C$ satisfy the conditions of Theorem 3. There exist constants $c_{1}$ and $c_{2}$ such that, for any modular point $(x_{1},x_{2},x_{3})\in C$,
[…]
Theorem 6.
Let $C$ and $(b_{1},b_{2},b_{3})$ be as in Theorem 4 and let $\epsilon>0$. There exist constants $c_{3}$, $c_{4}$ and $\delta$ such that, for any point modular point $s=(x_{1},x_{2},x_{3})\in C$ of supersingular exponent $\delta$,
[…]
We refer to [DO23, Section 1.D] for a detailed history of the G-function method for problems of unlikely intersections. As alluded to above, the results herein are a continuation of that method. The original motivation, and one of the key novelties, of this paper was to carry out this method in the interior of a Shimura variety, as opposed to at a point of degeneration at the boundary. Theorem 4, which is close in spirit to [And95, Théorème 1] (and whose proof is very much inspired by the proof of the latter) is our main result in this direction. This line of inquiry will be followed up in a series of forthcoming papers of the third author [Pap25a, Pap25b, Pap25c]
Structure of the paper
In Section 2, we give standard constructions pertaining to the G-functions method. In Section 3, we explain how to construct $v$-adic relations between the G-functions associated with elliptic schemes having isogenous fibres at a fixed point. In Section 4, we prove Theorems 5 and 6. In Section 5, we sketch the (now standard) arguments for deducing Theorems 3 and 4 from Theorems 5 and 6, respectively.
Acknowledgements
The first author is indebted to Gregorio Baldi, with whom he had several interesting discussions around the idea of intersections at the boundary. For part of this work, he was supported by a grant from the School of Mathematics at the Institute for Advanced Studies, Princeton.
The second author’s work on this project was supported by the Engineering and Physical Sciences Research Council (EP/Y020758/1).
The third author started work on this project while supported by Michael Temkin’s ERC Consolidator Grant BirNonArchGeom (770922). For the majority of the work, he received funding from the European Union (ERC, SharpOS, 101087910) and the ISRAEL SCIENCE FOUNDATION (grant No. 2067/23). Latterly, he was supported by the Minerva Research Foundation Member Fund, while in residence at the Institute for Advanced Study, for the academic year 2025-26.
The authors thank Jacob Tsimerman for kindly answering questions related to his work [BT25].
2. Preliminaries
In this section, we assemble various notation and background material, which will be used throughout the rest of the paper.
2.1. Matrices
For a square matrix $M$, we denote its trace by $\operatorname{tr}(M)$ and its determinant by $\det(M)$. We denote its adjugate by $M^{\rm adj}$ (which, if $M$ is invertible, is equal to $\det(M)M^{-1}$) and its transpose by $M^{T}$. For any ring $R$, we denote by $\mathrm{M}_{2}(R)$ the ring of $2\times 2$ matrices with coefficients in $R$.
2.2. Fields
For a number field $K$, we denote its ring of integers by $\mathcal{O}_{K}$ and its algebraic closure by $\overline{K}$. By a place $v$ of $K$, we refer to the equivalence class of an absolute value on $K$. We denote the set of places of $K$ by $\Sigma_{K}$, and we denote the set of archimedean (resp. finite, or non-archimedean) places of $K$ by $\Sigma_{K,\infty}$ (resp. $\Sigma_{K,f}$). For $v\in\Sigma_{K,f}$, we will use the following notations:
  • $K_{v}:=$ the completion of $K$ with respect to $v$;
  • $\mathcal{O}_{K_{v}}:=$ the ring of integers in $K_{v}$;
  • $\mathfrak{p}_{v}:=$ the prime ideal of $\mathcal{O}_{K}$ corresponding to $v$;
  • $k_{v}:=$ the residue field $\mathcal{O}_{K}/\mathfrak{p}_{v}$;
  • $W(k_{v}):=$ the ring of Witt vectors over $k_{v}$.
By abuse of notation, for $v\in\Sigma_{K,\infty}$, we define $K_{v}:=\mathbb{C}$ (even if the completion of $K$ with respect to $v$ is $\mathbb{R}$).
For $v\in\Sigma_{K}$, we denote by $|\cdot|_{v}$ the absolute value on $K_{v}$ extending the standard absolute value on $\mathbb{Q}$. We write $(\cdot)^{v\textrm{-}\mathrm{an}}$ for the analytification functor from schemes locally of finite type over $K_{v}$ to (rigid or complex) analytic spaces over $K_{v}$. For a power series $F\in K[[X]]$, we let $F^{v\textrm{-}\mathrm{an}}$ denote the corresponding analytic function defined on the open disc $D(0,R(F^{v\textrm{-}\mathrm{an}}),K_{v})$, where $R(F^{v\textrm{-}\mathrm{an}})$ is the $v$-adic radius of convergence.
2.3. Elliptic curves
For an $R$-scheme $X\to\operatorname{Spec}R$ and an $R$-algebra $R\to S$, we denote by $X_{S}$ the base-change $X\times_{\operatorname{Spec}R}\operatorname{Spec}S$. For an elliptic curve $E$ over $K$, we say that $E$ has good reduction at $v\in\Sigma_{K,f}$ if the special fibre $\mathcal{E}_{k_{v}}$ of the minimal Weierstrass model $\mathcal{E}$ of $E_{K_{v}}$ over $\mathcal{O}_{K_{v}}$ is non-singular. We say $E$ has bad reduction at $v$ otherwise. If $\mathcal{E}_{k_{v}}$ is non-singular, it is an elliptic curve over $k_{v}$, and we say that $E$ has ordinary (resp. supersingular) reduction at $v$ if $\mathcal{E}_{k_{v}}$ is ordinary (resp. supersingular). All of these properties are preserved by isogenies over $K$.
Now let $\mathcal{X}$ be a semiabelian scheme over $\operatorname{Spec}\mathcal{O}_{K}$ satisfying $\mathcal{X}_{K}\cong E$. Then, for any $v\in\Sigma_{K,f}$, the base-change $\mathcal{X}_{\mathcal{O}_{K_{v}}}$ is the connected Néron model of $E_{K_{v}}$ over $\operatorname{Spec}\mathcal{O}_{K_{v}}$ [BLR90, Prop. 7.4/3]. It follows that $\mathcal{X}_{k_{v}}$ is an elliptic curve if and only if $E$ has good reduction at $v$. Moreover, if $\mathcal{X}_{k_{v}}$ is an elliptic curve, it is ordinary (resp. supersingular) if and only if $E$ has ordinary (resp. supersingular) reduction at $v$.
2.4. Taylor series with respect to a local parameter
Let $C$ be a smooth geometrically irreducible algebraic curve defined over a number field $K$. Let $s_{0}\in C(K)$ and let $x\in K(C)$ be a local parameter at $s_{0}$ (which is to say that $x$ is a generator for the maximal ideal of the local ring $\mathcal{O}_{C,s_{0}}$). Let $v\in\Sigma_{K}$ and let $\mathfrak{m}$ denote the maximal ideal of the stalk $\mathcal{O}_{C,v,s_{0}}$ of $\mathcal{O}_{C^{v\textrm{-}\mathrm{an}}}$ at $s^{{v\textrm{-}\mathrm{an}}}_{0}$ (which is generated by the equivalence class of $x^{{v\textrm{-}\mathrm{an}}}$). Since $\mathcal{O}_{C,v,s_{0}}$ is Noetherian, it embeds into its $\mathfrak{m}$-adic completion $\hat{\mathcal{O}}_{C,v,s_{0}}$. Furthermore, since $\hat{\mathcal{O}}_{C,v,s_{0}}$ is regular, it is isomorphic to $K_{v}[[X]]$, and we may choose the isomorphism so that it sends the germ of $x^{v\textrm{-}\mathrm{an}}$ to the indeterminate $X$. Let $T$ denote the resulting injective ring homomorphism
$T:\mathcal{O}_{C,v,s_{0}}\to K_{v}[[X]].$
For any $f\in\mathcal{O}_{C,v,s_{0}}$, we refer to $T(f)$ as the Taylor series of $f$ with respect to $x$.
2.5. G-functions
A G-function is a power series $F(X)=\sum_{n\geq 0}a_{n}X^{n}$ with coefficients $a_{n}$ in a number field $K$ which satisfies the following conditions:
  1. (1)
[…]
  2. (2)
    there exists a sequence of positive integers $(d_{n})$, which grows at most geometrically, such that $d_{n}a_{m}$ is an algebraic integer for all $m\in\{1,\dotsc,n\}$;
  3. (3)
    $F$ satisfies a linear homogeneous differential equation
    $\frac{d^{\mu}}{dX^{\mu}}F+\gamma_{\mu-1}\frac{d^{\mu-1}}{dX^{\mu-1}}F+\dotsb+\gamma_{1}\frac{d}{dX}F+\gamma_{0}F=0$
    with coefficients $\gamma_{i}\in K(X)$.
2.6. G-functions associated with an abelian scheme
Consider the situation in Section 2.4 and let $\pi\colon\mathcal{A}\to C$ be an abelian scheme of relative dimension $g$, also defined over $K$. We write $H^{1}_{DR}(\mathcal{A}/C)$ for the relative algebraic de Rham cohomology, equipped with its Gauss–Manin connection
$\nabla:H^{1}_{DR}(\mathcal{A}/C)\to H^{1}_{DR}(\mathcal{A}/C)\otimes_{\mathcal{O}_{C}}\Omega^{1}_{C}.$
By [Del70, Prop. 6.14], the $\mathcal{O}_{C}$-sheaf $H^{1}_{DR}(\mathcal{A}/C)$ is locally-free. Suppose that it is free. Equivalently, we can choose a basis $\{\omega_{1},\ldots,\omega_{2g}\}$ of global sections. Now let $v\in\Sigma_{K}$ and let $\Delta_{v}$ be an open subspace of $C^{v\textrm{-}\mathrm{an}}$ containing $s^{v\textrm{-}\mathrm{an}}_{0}$ such that $H^{1}_{DR}(\mathcal{A}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ possesses a basis of horizontal sections $\{\gamma_{v,1},\ldots,\gamma_{v,2g}\}$. (For $v\in\Sigma_{K,\infty}$, the existence of $\Delta_{v}$ is classical—any simply connected open subset would suffice. For $v\in\Sigma_{K,f}$, the existence of $\Delta_{v}$ is guaranteed by [Ked22, Prop. 9.3.3].) Let
$\Omega_{v}\in\mathrm{M}_{2g}(\mathcal{O}_{C^{v\textrm{-}\mathrm{an}}}(\Delta_{v}))$
be the matrix whose rows give the coordinates of the $\omega^{{v\textrm{-}\mathrm{an}}}_{i}$ in terms of the $\gamma_{v,j}$.
Remark 7.
In the case $v\in\Sigma_{K,\infty}$, we may choose the horizontal sections $\gamma_{v,1},\dotsc,\gamma_{v,2g}$ so that they correspond to sections of the constant local system $R^{1}\pi^{v\textrm{-}\mathrm{an}}_{*}\overline{\mathbb{Q}}$ under the comparison isomorphism between de Rham and Betti cohomology. Then $\Omega_{v}$ is a matrix of complex periods of the family of abelian varieties $\mathcal{A}\to C$.
Lemma 8.
Define $Y_{v}:=\Omega_{v}\cdot\Omega_{v}(s^{v\textrm{-}\mathrm{an}}_{0})^{-1}$ and denote its entries by $Y_{v,mn}$ (that is, $1\leq m,n\leq 2g$). Let $F_{v,mn}$ denote the Taylor series of $Y_{v,mn}$ with respect to $x$. Then the $F_{v,mn}$ are G-functions, belonging to $K[[X]]$, independent of $v\in\Sigma_{K}$.
Proof.
It is straightforward to verify that the matrix $F_{v}:=(F_{v,mn})_{mn}$ satisfies a linear differential equation $\frac{d}{dx}(-)=A(-)$ induced from the action of $\nabla(\frac{d}{dx})$ on $H^{1}_{DR}(\mathcal{A}/C)$, where $A\in\mathrm{M}_{2g}(K((X)))$. Starting from the fact that $F_{v}(0)=I_{2g}$ has entries in $K$, it follows that $F_{v,nm}\in K[[X]]$ for all $m,n\in\{1,\ldots,2g\}$. By [And89, p3, Thm. B], therefore, $F_{v,nm}$ is a G-function for all $m,n\in\{1,\ldots,2g\}$. Furthermore, $F_{v}$ is the unique solution of $\frac{d}{dx}(-)=A(-)$ in $\mathrm{M}_{2g}(K((X)))$ satisfying $F_{v}(0)=I_{2g}$, so it is independent of $v$. ∎
By virtue of Lemma 8, we denote $F_{v,mn}$ by $F_{mn}$ and we refer to the $F_{mn}$ for $1\leq m,n\leq 2g$ as the G-functions associated with $\mathcal{A}\to C$, $x$, and $\{\omega_{1},\ldots,\omega_{2g}\}$.
2.7. Relations
Let $F_{1},\dotsc,F_{n}\in\overline{\mathbb{Q}}[[X]]$. We say that a homogeneous polynomial $\tilde{Q}\in\overline{\mathbb{Q}}[X][X_{1},\dotsc,X_{n}]$ is a functional relation between $F_{1},\dotsc,F_{n}$ if
$\tilde{Q}(X)(F_{1}(X),\dotsc,F_{n}(X))=0\text{ in }\overline{\mathbb{Q}}[[X]].$
Let $\xi\in\overline{\mathbb{Q}}$ and let $Q\in\overline{\mathbb{Q}}[X_{1},\dotsc,X_{n}]$ be a homogeneous polynomial. Let $K$ be a number field which contains $\xi$ and all of the coefficients of $F_{1},\dotsc,F_{n}$ and $Q$. We use the following definitions.
  1. (1)
    $Q$ is an $v$-adic relation between the evaluations at $\xi$ of $F_{1},\dotsc,F_{n}$ (for $v\in\Sigma_{K}$) if $\lvert\xi\rvert_{v}<\min\{1,R(F_{1}^{v\textrm{-}\mathrm{an}}),\dotsc,R(F_{n}^{v\textrm{-}\mathrm{an}})\}$ and
    $Q^{v\textrm{-}\mathrm{an}}(F_{1}^{v\textrm{-}\mathrm{an}}(\xi^{v\textrm{-}\mathrm{an}}),\dotsc,F_{n}^{v\textrm{-}\mathrm{an}}(\xi^{v\textrm{-}\mathrm{an}}))=0.$
  2. (2)
    $Q$ is a global relation between the evaluations at $\xi$ of $F_{1},\dotsc,F_{n}$ if it is a $v$-adic relation between the evaluations at $\xi$ for every place $v$ of $K$ satisfying $\lvert\xi\rvert_{v}<\min\{1,R(F_{1}^{v\textrm{-}\mathrm{an}}),\dotsc,R(F_{n}^{v\textrm{-}\mathrm{an}})\}$.
  3. (3)
    $Q$ is a trivial relation between $F_{1},\dotsc,F_{n}$ at $\xi$ if it is the specialisation at $X=\xi$ of a functional relation $\tilde{Q}\in\overline{\mathbb{Q}}[X][X_{1},\dotsc,X_{n}]$ between $F_{1},\dotsc,F_{n}$, where $\tilde{Q}$ is homogeneous of the same degree as $Q$.
2.8. Simple neighbourhoods
Return to the situation of Section 2.4 and let $\mathfrak{C}$ be a regular integral flat $\mathcal{O}_{K}$-scheme with an isomorphism $\mathfrak{C}_{K}\to C$. Let $v\in\Sigma_{K,f}$ and let $\mathfrak{C}_{v{\dagger}}$ denote the formal completion of $\mathfrak{C}$ along the fibre $\mathfrak{C}_{k_{v}}$. Let $\mathfrak{C}_{v{\dagger}}^{\mathrm{rig}}$ denote the associated rigid analytic space (see [Ber96, 0.2.6]). Then there is a canonical open immersion $\mathfrak{C}_{v{\dagger}}^{\mathrm{rig}}\to C^{v\textrm{-}\mathrm{an}}$ [Ber96, Prop. 0.3.5]. We say that an open subspace $U$ of $C^{v\textrm{-}\mathrm{an}}$ is simple if $U\subset\mathfrak{C}_{v{\dagger}}^{\mathrm{rig}}$ and the reduction map $\operatorname{red}_{v}:\mathfrak{C}_{v{\dagger}}^{\mathrm{rig}}\to\mathfrak{C}_{k_{v}}$ is constant on $U$.
The following is standard.
Lemma 9.
Let $U$ be a simple open subspace of $C^{v\textrm{-}\mathrm{an}}$ and let $t\in\mathfrak{C}(k_{v})$ denote the (constant) value of $\operatorname{red}_{v}$ on $U$. Then $\operatorname{red}_{v}^{-1}(t)$ is equal to the completion $\mathfrak{C}_{v,\mathrm{for}{\dagger}}$ of $\mathfrak{C}$ along $t$. If $t$ is a smooth point of $\mathfrak{C}_{k_{v}}$, then $\operatorname{red}_{v}^{-1}(t)$ is isomorphic to the rigid open unit disc.
3. Relations for elliptic schemes
In this section, we construct $v$-adic relations between the G-functions associated with elliptic schemes having isogenous fibres over a fixed base point. In Section 3.1, we consider a pair of elliptic schemes. In Section 3.3, we introduce a third elliptic scheme, in order to obtain a single relation for all places of good ordinary reduction of the fixed fibres.
3.1. Isogenous pairs
Let $C^{\prime}$ be a smooth projective geometrically irreducible algebraic curve, defined over a number field $K$, and let $\mathfrak{C}^{\prime}$ be a regular $\mathcal{O}_{K}$-model of $C^{\prime}$ (see [Liu06, Definition 10.1.1] for the definition; such a model exists by [Liu06, Cor. 8.3.45]).
Definition.
[…]
For $i\in\{1,2\}$, let $\mathfrak{G}^{\prime}_{i}\to\mathfrak{C}^{\prime}$ be a semiabelian scheme such that for some non-empty Zariski open subset $C\subset C^{\prime}$ the base-change $\mathcal{E}_{i}:=\mathfrak{G}^{\prime}_{i}\times_{\mathfrak{C}^{\prime}}C$ is an elliptic scheme over $C$. For $i\in\{1,2\}$, let $\mathfrak{C}\subset\mathfrak{C}^{\prime}$ denote a non-empty Zariski open subscheme such that $\mathfrak{G}_{i}:=\mathfrak{G}^{\prime}_{i}\times_{\mathfrak{C}^{\prime}}\mathfrak{C}$ is an elliptic scheme over $\mathfrak{C}$ (such a subscheme exists because the toric rank is upper semi-continuous [Lan13, Lemma 3.3.1.4]). Suppose that $\mathfrak{C}_{K}=C$.
Let $s_{0}\in C(K)$ and suppose that there exists a $K$-isogeny $f_{0}:\mathcal{E}_{1,s_{0}}\to\mathcal{E}_{2,s_{0}}$ of degree $N_{0}:=\deg(f_{0})$. We will say that $v\in\Sigma_{K,f}$ is a place of good (resp. bad) reduction if one (equivalently, all) of the $\mathcal{E}_{i,s_{0}}$ have good (resp. bad) reduction at $v$.
Choose bases $\{\omega_{i},\eta_{i}\}$ for $H^{1}_{DR}(\mathcal{E}_{i}/C)$ such that
(1) $\omega_{1,s_{0}}=f^{*}_{0}\omega_{2,s_{0}}\text{ and }\eta_{1,s_{0}}=f^{*}_{0}\eta_{2,s_{0}},$
where
$f^{*}_{0}:H^{1}_{DR}(\mathcal{E}_{2,s_{0}}/K)\to H^{1}_{DR}(\mathcal{E}_{1,s_{0}}/K)$
denotes the pullback induced by $f_{0}$. Let $x\in K(C)$ be a local parameter at $s_{0}$ and let $F_{imn}\in K[[X]]$ denote the G-functions associated with $\mathcal{E}_{i}\to C$, $x$, and $\{\omega_{i},\eta_{i}\}$. Let
$\mathcal{G}:=\{F_{imn}:1\leq i,m,n\leq 2\},$
and, for each $v\in\Sigma_{K}$, define
$R_{v}:=\min\{1,R(F^{v\textrm{-}\mathrm{an}}):F\in\mathcal{G}\}.$
For each $v\in\Sigma_{K,f}$, let $r_{v}$ and $U_{v}$ denote, respectively, the positive real number and the open subspace of $C^{v\textrm{-}\mathrm{an}}$ afforded to us by Proposition 20. For each $v\in\Sigma_{K,\infty}$, let $U_{v}$ denote an open subspace of $C^{v\textrm{-}\mathrm{an}}$ such that $x^{v\textrm{-}\mathrm{an}}$ restricts to a biholomorphism from $U_{v}$ to $D(0,r_{v},K_{v})$ for some $r_{v}>0$.
For each $v\in\Sigma_{K}$, define
$R^{\leq}_{v}:=\min\{R_{v},r_{v}\}\leq r_{v}$
and $\Delta_{v}:=(x^{v\textrm{-}\mathrm{an}}|_{U_{v}})^{-1}(D(0,R^{\leq}_{v},K_{v}))$. By shrinking the $r_{v}$ (if necessary), we can and do assume that $\Delta_{v}$ is
  • (a)
    simply connected for $v\in\Sigma_{v,\infty}$;
  • (b)
    contained in the open subspace $\mathcal{C}_{v}$ of Section 3.2.2 for the (finitely many) $v\in\Sigma_{v,f}$ of bad reduction.
By Lemma 9, we can and do also suppose that $\Delta_{v}$ is
  • (c)
    simple for all $v\in\Sigma_{K,f}$ of good reduction.
Lemma 10.
[…]
$\mathfrak{G}^{\prime}_{1,\mathfrak{s},k_{\hat{v}}}\cong\mathfrak{G}^{\prime}_{1,\mathfrak{s}_{0},k_{\hat{v}}}\text{ and }\mathfrak{G}^{\prime}_{2,\mathfrak{s},k_{\hat{v}}}\cong\mathfrak{G}^{\prime}_{2,\mathfrak{s}_{0},k_{\hat{v}}}$.
Proof.
[…]
Now we state the main result of this section.
Proposition 11.
[…]
Let $v\in\Sigma_{K}$ and choose $\hat{v}\in\Sigma_{\hat{K}}$ lying over $v$.
If $s^{\hat{v}\textrm{-}\mathrm{an}}\in\Delta_{v}$, there exists a $\hat{v}-$adic relation $Q_{\hat{v}}$ of degree at most $2$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$ which is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i,m,n\leq 2]$ generated by the elements
$\det((X_{1mn})_{mn})-\det((X_{2mn})_{mn}).$
3.2. Proof of Proposition 11
If necessary, replace $K$ with $\hat{K}$ and $v$ with $\hat{v}$, and, for uniformity of notation, define the field
$F_{v}:=\begin{cases}\mathbb{Q}&\quad\text{if }v\in\Sigma_{K,\infty}\\ K_{v}&\quad\text{if }v\in\Sigma_{K,f}.\end{cases}$
(This is the field of coefficients for the “$v$-adic cohomology” theory i.e. Betti cohomology when $v\in\Sigma_{K,\infty}$ or crystalline cohomology when $v\in\Sigma_{K,f}$.) We will split the proof of Proposition 3.2 into two cases.
3.2.1. Case 1
Suppose $v\in\Sigma_{K,\infty}$ or that $v\in\Sigma_{K,f}$ and the $\mathcal{E}_{i,s}$ have good reduction at $v$ (which is to say that the $\mathcal{E}_{i,s,k_{v}}=\mathfrak{G}^{\prime}_{i,\mathfrak{s},k_{v}}$ are elliptic curves). Define the $F_{v}$-vector space
$H^{1}_{v}(\mathcal{E}_{i,s}):=\begin{cases}H^{1}(\mathcal{E}_{i,s}^{v\textrm{-}\mathrm{an}},F_{v})&\quad\text{if }v\in\Sigma_{K,\infty}\\ H^{1}_{cris}(\mathcal{E}_{i,s,k_{v}}/W(k_{v}))\otimes_{W(k_{v})}F_{v}&\quad\text{if }v\in\Sigma_{K,f},\end{cases}$
where the former denotes singular cohomology with $\mathbb{Q}$-coefficients, and the latter denotes crystalline cohomology with $W(k_{v})$-coefficients [Ber74] tensored with $F_{v}$. Let $H^{1}_{DR}(\mathcal{E}_{i,s}/K_{v})$ denote the de Rham cohomology of $\mathcal{E}_{i,s,K_{v}}$. Note that $H^{1}_{DR}(\mathcal{E}_{i,s}/K_{v})\cong H^{1}_{DR}(\mathcal{E}_{i,s}^{v\textrm{-}\mathrm{an}})$ ([ABC20, Cor. 31.1.2, 32.2.2]). Define the analogous objects for $s$ replaced with $s_{0}$. For $i\in\{1,2\}$, let $\pi_{i}:\mathcal{E}_{i}\to C$ denote the structure map.
Lemma 12.
For $t\in\{s,s_{0}\}$, there exist canonical (comparison) isomorphisms
$\sigma_{i}:H^{1}_{DR}(\mathcal{E}_{i,t}/K_{v})\to H^{1}_{v}(\mathcal{E}_{i,t})\otimes_{F_{v}}K_{v}$
and commutative diagrams
$H^{1}_{DR}(\mathcal{E}_{i,s_{0}}/K_{v})$$H^{1}_{DR}(\mathcal{E}_{i,s}/K_{v})$$H^{1}_{v}(\mathcal{E}_{i,s_{0}})\otimes_{F_{v}}K_{v}$$H^{1}_{v}(\mathcal{E}_{i,s})\otimes_{F_{v}}K_{v}$$\scriptstyle{\sigma_{i}}$$\scriptstyle{\epsilon_{i}}$$\scriptstyle{\sigma_{i}}$$\scriptstyle{\iota_{i}}$
where $\epsilon_{i}$ is the isomorphism induced by the Gauss–Manin connection (“parallel transport”) and $\iota_{i}$ is induced
(if $v\in\Sigma_{K,\infty}$) by the canonical isomorphism $R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{s_{0}}\cong R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{s}$ of fibres of the constant sheaf $R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{\Delta_{v}}$.;
(if $v\in\Sigma_{K,f}$) by the pullback $\alpha^{*}_{i}$ associated with an isogeny $\alpha_{i}:\mathcal{E}_{i,s,k_{v}}\to\mathcal{E}_{i,s_{0},k_{v}}$ whose degree we note $d_{i}$.
Proof.
If $v\in\Sigma_{K,\infty}$, the maps $\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}|_{\Delta_{v}}\to\Delta_{v}$ are smooth and proper, and so the claims follow from [Kat72, Prop. 4.1.2]. For $v\in\Sigma_{K,f}$, the $\sigma_{i}$ are afforded to us by [BO83, (2.4.2)]. Furthermore, in the terminology of [Ogu84, §5], $\mathcal{E}_{i,K_{v}}$ has good reduction over $\mathfrak{C}_{\mathcal{O}_{K_{v}}}$. Therefore, the commutative diagrams in this case are afforded to us by [Ogu84, Rem. 5.14.3] and Lemma 10. ∎
Now we construct relations. To that end, let $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ denote a $K_{v}$-basis for $H^{1}_{DR}(\mathcal{E}_{2,s_{0}}/K_{v})$ whose elements are contained in $\sigma^{-1}_{2}(H^{1}_{v}(\mathcal{E}_{2,s_{0}}))$, and let
(2) $\gamma_{v,1,s_{0}}:=f^{*}_{0}\gamma_{v,2,s_{0}},\ \delta_{v,1,s_{0}}:=f^{*}_{0}\delta_{v,2,s_{0}}.$
Since $f_{0}$ is an isogeny, $\gamma_{v,1,s_{0}}$ and $\delta_{v,2,s_{0}}$ form a $K_{v}$-basis for $H^{1}_{DR}(\mathcal{E}_{1,s_{0}}/K_{v})$.
We claim there is a commutative diagram
$H^{1}_{DR}(\mathcal{E}_{2,s_{0}}/K_{v})$$H^{1}_{DR}(\mathcal{E}_{1,s_{0}}/K_{v})$$H^{1}_{v}(\mathcal{E}_{2,s_{0}})\otimes_{F_{v}}K_{v}$$H^{1}_{v}(\mathcal{E}_{1,s_{0}})\otimes_{F_{v}}K_{v}$$\scriptstyle{\sigma_{2}}$$\scriptstyle{f^{*}_{0}}$$\scriptstyle{\sigma_{1}}$$\scriptstyle{f^{*}_{0}}$
From this, we conclude $\gamma_{v,1,s_{0}},\delta_{v,1,s_{0}}\in\sigma^{-1}_{1}(H^{1}_{v}(\mathcal{E}_{1,s_{0}}))$. For $v\in\Sigma_{K,\infty}$, the existence of the diagram is standard. For $v\in\Sigma_{K,f}$, we consider $\mathfrak{t}\in\{\mathfrak{s}_{0},\mathfrak{s}\}$ and write $t:=\mathfrak{t}_{K}$. Then, the semiabelian scheme $\mathfrak{G}_{i,\mathfrak{t}}$ is the connected Néron model of its generic fibre $\mathcal{E}_{i,t}$ and so, by the Néron mapping property, we obtain an isogeny
$f_{0,k_{v}}:\mathcal{E}_{1,s_{0},k_{v}}\to\mathcal{E}_{2,s_{0},k_{v}}$
(and similarly $f_{k_{v}}:\mathcal{E}_{1,s,k_{v}}\to\mathcal{E}_{2,s,k_{v}}$). This proves existence of the diagram. The fact that the diagram commutes in this case is [BO83, (2.4.5)].
The isomorphisms $\epsilon_{i}$ allow us to extend the $\{\gamma_{v,i,s_{0}},\delta_{v,i,s_{0}}\}$ to horizontal bases $\{\gamma_{v,i},\delta_{v,i}\}$ for $H^{1}_{DR}(\mathcal{E}_{i}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ and we let $\Omega_{v,i}\in\mathrm{M}_{2}(\mathcal{O}_{\Delta_{v}})$ be the matrix whose rows give the coordinates of $\{\omega^{v\textrm{-}\mathrm{an}}_{i},\eta^{v\textrm{-}\mathrm{an}}_{i}\}$ in terms of $\{\gamma_{v,i},\delta_{v,i}\}$. By (1) and (2), we have $\Omega_{v,0}:=\Omega_{v,1}(s^{v\textrm{-}\mathrm{an}}_{0})=\Omega_{v,2}(s^{v\textrm{-}\mathrm{an}}_{0})$ and we define $Y_{v,i}:=\Omega_{v,i}\cdot\Omega_{v,0}^{-1}$. By construction, the Taylor series of the entries of the $Y_{v,i}$ with respect to $x$, are precisely the $F_{imn}$ invoked in Section 3.1.
Write
$f^{*}\omega_{2,s}=a\omega_{1,s}+c\eta_{1,s}\text{ and }f^{*}\eta_{2,s}=b\omega_{1,s}+d\eta_{1,s}$
for some $a,b,c,d\in K$, and
$f^{{v\textrm{-}\mathrm{an}}*}\gamma_{v,2,s}=p_{v}\gamma_{v,1,s}+q_{v}\delta_{v,1,s}\text{ and }f^{{v\textrm{-}\mathrm{an}}*}\delta_{v,2,s}=r_{v}\gamma_{v,1,s}+s_{v}\delta_{v,1,s}$
for some $p_{v},q_{v},r_{v},s_{v}\in F_{v}$. We consider these coefficients as forming matrices
$A:=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\text{ and }M_{v}:=\left(\begin{matrix}p_{v}&r_{v}\\ q_{v}&s_{v}\end{matrix}\right).$
Lemma 13.
We have $\operatorname{tr}(M_{v})\in\mathbb{Q}$.
Proof.
Again, we have a commutative diagram
$H^{1}_{DR}(\mathcal{E}_{2,s}/K_{v})$$H^{1}_{DR}(\mathcal{E}_{1,s}/K_{v})$$H^{1}_{v}(\mathcal{E}_{2,s})\otimes_{F_{v}}K_{v}$$H^{1}_{v}(\mathcal{E}_{1,s})\otimes_{F_{v}}K_{v}$$\scriptstyle{\sigma_{2}}$$\scriptstyle{f^{*}}$$\scriptstyle{\sigma_{1}}$$\scriptstyle{f^{*}}$
If $v\in\Sigma_{K,\infty}$, the bottom arrow is the extension to $K_{v}$ of an isomorphism $H^{1}_{v}(\mathcal{E}_{2,s})\to H^{1}_{v}(\mathcal{E}_{1,s})$ of $\mathbb{Q}$-vector spaces. By Lemma 12, the sets
$\{\sigma_{2}(\gamma_{v,2,s}),\sigma_{2}(\delta_{v,2,s)}\}\text{ and }\{\sigma_{1}(\gamma_{v,1,s}),\sigma_{1}(\delta_{v,2,s})\}$
are bases $B_{2}$ and $B_{1}$ for $H^{1}_{v}(\mathcal{E}_{2,s})$ and $H^{1}_{v}(\mathcal{E}_{1,s})$, respectively. The matrix representing the lower $f^{*}$ with respect to $B_{2}$ and $B_{1}$ is $M_{v}$, and so $M_{v}\in\mathrm{M}_{2}(\mathbb{Q})$. In particular, $\operatorname{tr}(M_{v})\in\mathbb{Q}$, as claimed.
For $v\in\Sigma_{K,f}$, we write $f^{\vee}_{0,k_{v}}:\mathcal{E}_{2,s_{0},k_{v}}\to\mathcal{E}_{1,s_{0},k_{v}}$ for the dual isogeny to $f_{0,k_{v}}$. Then, the composition
$\mathcal{E}_{2,s_{0},k_{v}}$$\mathcal{E}_{1,s_{0},k_{v}}$$\mathcal{E}_{1,s,k_{v}}$$\mathcal{E}_{2,s,k_{v}}$$\mathcal{E}_{2,s_{0},k_{v}}$$\scriptstyle{f^{\vee}_{0,k_{v}}}$$\scriptstyle{\alpha_{1}^{\vee}}$$\scriptstyle{f_{k_{v}}}$$\scriptstyle{\alpha_{2}}$
yields an endomorphism of $\mathcal{E}_{2,s_{0},k_{v}}$ and hence an endomorphism of $H^{1}_{v}(\mathcal{E}_{2,s_{0}})$, which we denote $\phi$. By Lemma 12, this forms part of the commutative diagram
$H^{1}_{DR}(\mathcal{E}_{2,s_{0}})$$H^{1}_{DR}(\mathcal{E}_{2,s})$$H^{1}_{DR}(\mathcal{E}_{1,s})$$H^{1}_{DR}(\mathcal{E}_{1,s_{0}})$$H^{1}_{DR}(\mathcal{E}_{2,s_{0}})$$H^{1}_{v}(\mathcal{E}_{2,s_{0}})$$H^{1}_{v}(\mathcal{E}_{2,s})$$H^{1}_{v}(\mathcal{E}_{1,s})$$H^{1}_{v}(\mathcal{E}_{1,s_{0}})$$H^{1}_{v}(\mathcal{E}_{2,s_{0}}).$$\scriptstyle{\sigma_{2}}$$\scriptstyle{\epsilon_{2}}$$\scriptstyle{\sigma_{2}}$$\scriptstyle{f^{*}}$$\scriptstyle{\sigma_{1}}$$\scriptstyle{d_{1}\epsilon_{1}^{-1}}$$\scriptstyle{\sigma_{1}}$$\scriptstyle{N_{0}(f_{0}^{*})^{-1}}$$\scriptstyle{\sigma_{2}}$$\scriptstyle{\iota_{2}}$$\scriptstyle{\phi}$$\scriptstyle{f^{*}}$$\scriptstyle{d_{1}\iota_{1}^{-1}}$$\scriptstyle{N_{0}(f_{0}^{*})^{-1}}$
By following $\gamma_{v,2,s_{0}}$ and $\delta_{v,2,s_{0}}$ along the top line of this diagram, we conclude that, with respect to the basis $\{\sigma_{2}(\gamma_{v,2,s_{0}}),\sigma_{2}(\delta_{v,2,s_{0}})\}$, the matrix representing $\phi$ is $d_{1}N_{0}M_{v}$.
Therefore, since $H^{1}_{cris}(\mathcal{E}_{2,s_{0},k_{v}}/W(k_{v}))$ is canonically isomorphic to the Dieudonné module associated with the $p$-divisible group $\mathcal{E}_{2,s_{0},k_{v}}[p^{\infty}]$ [BC09, Rem. 7.3.3], we conclude from [Dem86, V.2, Corollary] that $\operatorname{tr}(d_{1}N_{0}M_{v})\in\mathbb{Z}$. We obtain $\operatorname{tr}(M_{v})\in\mathbb{Q}$, as claimed. ∎
Resuming from the passage preceding Lemma 13, we have
$A^{T}\Omega_{v,1}(s^{v\textrm{-}\mathrm{an}})=\Omega_{v,2}(s^{v\textrm{-}\mathrm{an}})M^{T}_{v}.$
Since the Taylor series of an analytic function $\varphi$ converges to $\varphi$ inside the radius of convergence, if we write $F_{i,s}:=(F^{v\textrm{-}\mathrm{an}}_{i}(x^{v\textrm{-}\mathrm{an}}(s^{v\textrm{-}\mathrm{an}}))_{mn})_{1\leq m,n\leq 2}$, then we obtain
(3) $(F_{2,s})^{-1}A^{T}F_{1,s}=\Omega_{v,0}M^{T}_{v}\Omega_{v,0}^{-1}.$
Using the invariance of trace under conjugation, this yields
$\operatorname{tr}((F_{2,s})^{-1}A^{T}F_{1,s})=\operatorname{tr}(M^{T}_{v})=\operatorname{tr}(M_{v}),$
and, finally, multiplying both sides by $\det(F_{2,s})$, we arrive at
$\operatorname{tr}(F_{2,s}^{\rm adj}A^{T}F_{1,s})=\det(F_{2,s})\operatorname{tr}(M_{v}).$
This gives a $v$-adic relation $Q_{v}$ of degree $2$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$. The fact that $Q_{v}\notin I$ is readily seen by observing that $Q_{v}$ has degree $1$ with respect to the entries of $F_{1,s}$, while $I$ is generated by a polynomial of degree $2$ with respect to $X_{1mn}$ ($1\leq m,n\leq 2$).
3.2.2. Case 2
Now suppose $v\in\Sigma_{K,f}$ and that the $\mathcal{E}_{i,s}$ (equivalently, the $\mathcal{E}_{i,s_{0}}$) have bad reduction at $v$. Thus, after possibly replacing $K$ with a finite extension, we have $\mathfrak{G}^{\prime}_{i,\mathfrak{s}_{0},k_{v}}\cong\mathfrak{G}^{\prime}_{i,\mathfrak{s},k_{v}}\cong\mathbb{G}_{m,k_{v}}$. This implies that we can put ourselves in the situation of [DO23, 3.B.1] using the following dictionary, which sends the notations of [DO23, 3.B.1] to those of this section:
$R\mapsto\ \mathcal{O}_{K_{v}},\ \mathfrak{C}\mapsto\ \mathfrak{C}^{\prime}_{\mathcal{O}_{K_{v}}},\ \mathfrak{C}_{0,\mathfrak{p}}\mapsto\ \mathfrak{s}_{0}({\operatorname{Spec}{k_{v}}}),\ \mathfrak{G}\mapsto$ $\ \mathfrak{G}^{\prime}_{\mathcal{O}_{K_{v}}.}$
By [DO23, Lemma 3.3], for some open subspace $\mathcal{C}_{v}$ of $C^{{v\textrm{-}\mathrm{an}}}$ (which was referred to in Section 3.1 (c)) and $\mathcal{E}_{v,i}:=\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}|_{\mathcal{C}_{v}}$, there exist rigid analytic uniformisations $\phi_{v,i}:\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m}\times\mathcal{C}_{v}\to\mathcal{E}_{v,i}$. Moreover, by [Ger70, Satz 5], we obtain commutative diagrams
${\mathbb{G}_{m}^{v\textrm{-}\mathrm{an}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{v,1,s_{0}}}$$\scriptstyle{[m_{0}]}$${\mathbb{G}_{m}^{v\textrm{-}\mathrm{an}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{v,2,s_{0}}}$${\mathcal{E}_{v,1,s_{0}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{v\textrm{-}\mathrm{an}}_{0}}$${\mathcal{E}_{v,2,s_{0}}}$    ${\mathbb{G}_{m}^{v\textrm{-}\mathrm{an}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{v,1,s}}$$\scriptstyle{[m]}$${\mathbb{G}_{m}^{v\textrm{-}\mathrm{an}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{v,2,s}}$${\mathcal{E}_{v,1,s}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{v\textrm{-}\mathrm{an}}}$${\mathcal{E}_{v,2,s}}$
for natural numbers $m_{0}|N_{0}$ and $m|\deg(f)$, from which we obtain commutative diagrams
${H^{1}_{DR}(\mathcal{E}_{v,2,s_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,2,s_{0}}}$$\scriptstyle{(f^{v\textrm{-}\mathrm{an}}_{0})^{*}}$${H^{1}_{DR}(\mathcal{E}_{v,1,s_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,1,s_{0}}}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[m_{0}]}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})}$    ${H^{1}_{DR}(\mathcal{E}_{v,2,s})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,2,s}}$$\scriptstyle{(f^{v\textrm{-}\mathrm{an}})^{*}}$${H^{1}_{DR}(\mathcal{E}_{v,1,s})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,1,s}}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[m]}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m}).}$
Consider the natural injective map
$H^{1}_{DR}(\mathcal{E}_{i}/C)^{v\textrm{-}\mathrm{an}}\to H^{1}_{DR}(\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}/C^{v\textrm{-}\mathrm{an}})$
of modules with integrable connection. We also have the pullback
$\phi_{v,i}^{*}|_{\mathcal{C}_{v}}:H^{1}_{DR}(\mathcal{E}_{v,i}/\mathcal{C}_{v})\to H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m}\times\mathcal{C}_{v}/\mathcal{C}_{v})$
(which is non-trivial because $\phi_{v,i}$ is smooth), for which the target is isomorphic to
$H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})\otimes_{K_{v}}\mathcal{O}_{\mathcal{C}_{v}}=\mathcal{O}_{\mathcal{C}_{v}}\cdot dz/z.$
(Indeed, $\mathbb{G}_{m}\times C\to C$ is a rational elementary fibration [ABC20, Def. 25.1.4], hence, its Artin set is equal to $C$ [ABC20, Rmk. 25.2.3]. Therefore, [ABC20, Thm. 32.2.1] gives us this isomorphism.) Since $\Delta_{v}$ is contained in $\mathcal{C}_{v}$, it follows that we can choose a horizontal basis $\{\gamma_{v,2},\delta_{v,2}\}$ for $H^{1}_{DR}(\mathcal{E}_{2}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ such that, under the above morphisms, we have
$\gamma_{v,2}\mapsto dz/z,\ \delta_{v,2}\mapsto 0.$
As in Section 3.2.1, define the basis
$\gamma_{v,1,s_{0}}:=(f^{{v\textrm{-}\mathrm{an}}}_{0})^{*}\gamma_{v,1,s_{0}},\ \delta_{v,1,s_{0}}:=(f^{{v\textrm{-}\mathrm{an}}}_{0})^{*}\delta_{v,1,s_{0}}$
for $H^{1}_{DR}(\mathcal{E}_{v,1,s_{0}})$ and extend to a horizontal basis $\{\gamma_{v,1},\delta_{v,1}\}$ for $H^{1}_{DR}(\mathcal{E}_{2}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$. For the above morphisms, we have
$\gamma_{v,1}\mapsto m_{0}\,dz/z,\ \delta_{v,1}\mapsto 0,$
and, with $A$ and $\Omega_{v,i}$ defined as in Section 3.2.1, we arrive at the following equation:
$m_{0}\,A^{T}\Omega_{v,1}(s^{v\textrm{-}\mathrm{an}})(dz/z,0)^{T}=m\,\Omega_{v,2}(s^{v\textrm{-}\mathrm{an}})(dz/z,0)^{T}$
(where we write $(dz/z,0)^{T}$ for the corresponding column vector). If $(\Omega_{v,0})_{21}\neq 0$, we write $\lambda_{v,0}:=(\Omega_{v,0})_{11}(\Omega_{v,0})^{-1}_{21}$ and, expanding, we obtain
(4) $a(F^{v\textrm{-}\mathrm{an}}_{1,s})_{12}+c(F^{v\textrm{-}\mathrm{an}}_{1,s})_{22}-\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{12}$ $=\lambda_{v,0}[\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{11}-a(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-c(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}]$
(5) $b(F^{v\textrm{-}\mathrm{an}}_{1,s})_{12}+d(F^{v\textrm{-}\mathrm{an}}_{1,s})_{22}-\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{22}$ $=\lambda_{v,0}[\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{21}-b(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-d(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}],$
where, again, $(F^{v\textrm{-}\mathrm{an}}_{i,s})_{jk}:=(F^{v\textrm{-}\mathrm{an}}_{i}(x^{v\textrm{-}\mathrm{an}}(s^{v\textrm{-}\mathrm{an}})))_{jk}$. If the right-hand side of (4) is 0, we obtain a linear $v$-adic relation between the evaluations of the elements of $\mathcal{G}$ at $x(s)$. Otherwise, we can divide (5) by (4) and clear denominators to obtain a $v$-adic relation of degree $2$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$. If $(\Omega_{v,0})_{21}=0$, we must have $(\Omega_{v,0})_{11}\neq 0$, and we obtain
(6) $\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{11}-a(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-c(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}=0$
Again, the fact that any of these relations is not in $I$ is readily seen by elementary computation. ∎
3.3. Isogenous triples
In this section, we introduce a third elliptic scheme in order to obtain a single relation holding at all places of good ordinary reduction. That is, we repeat the setup of Section 3.1 but with the following modifications: we let $i\in\{1,2,3\}$, we suppose that there are isogenies
$f_{0}:\mathcal{E}_{1,s_{0}}\to\mathcal{E}_{2,s_{0}}\text{ and }f^{\prime}_{0}:\mathcal{E}_{1,s_{0}}\to\mathcal{E}_{3,s_{0}}$
of degrees $N_{0}:=\deg(f_{0})$ and $N^{\prime}_{0}:=\deg(f^{\prime}_{0})$, and we impose the analogous conditions on the bases for each $H^{1}_{DR}(\mathcal{E}_{i}/C)$. We will use the notations of Section 3.2.1.
The result is as follows.
Proposition 14.
[…]
$f:\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}\text{ and }f^{\prime}:\mathcal{E}_{1,s}\to\mathcal{E}_{3,s}.$
Let $v\in\Sigma_{K}$ and choose $\hat{v}\in\Sigma_{\hat{K}}$ lying over $v$.
If $s^{\hat{v}\textrm{-}\mathrm{an}}\in\Delta_{v}$, there exists a $\hat{v}-$adic relation $Q_{\hat{v}}$ of degree at most $4$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$ which is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i\leq 3,\ 1\leq m,n\leq 2]$ generated by the elements
$\det((X_{1mn})_{mn})-\det((X_{2mn})_{mn})\text{ and }\det((X_{1mn})_{mn})-\det((X_{3mn})_{mn}).$
Moreover, for $\hat{v}\in\Sigma_{\hat{K},f}$ for which the $\mathcal{E}_{i,s,k_{v}}$ are ordinary, we can choose $Q_{\hat{v}}$ independently of $v$.
Proof.
The first part of the proposition follows immediately from Proposition 11 (we may ignore $\mathcal{E}_{3}$ for this part).
To prove the second part, we replace $K$ with $\hat{K}$ and $v$ with $\hat{v}$, and we suppose that $v\in\Sigma_{K,f}$ is such that the $\mathcal{E}_{i,s,k_{v}}$ are ordinary.
In this case, $L:=\operatorname{End}(\mathcal{E}_{2,s,k_{v}})\otimes_{\mathbb{Z}}\mathbb{Q}$ is an imaginary quadratic field. After possibly replacing $K$ with a finite extension (which may depend on $v$ but will not be visible in our sought relation), we may choose $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ so that the action of $L$ on $H^{1}_{v}(\mathcal{E}_{2,s})$ is diagonal. This $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ determines $\{\gamma_{v,3,s_{0}},\delta_{v,3,s_{0}}\}$ and, with respect to its image in $H^{1}_{v}(\mathcal{E}_{3,s})$, the action of $\operatorname{End}(\mathcal{E}_{3,s,k_{v}})\otimes_{\mathbb{Z}}\mathbb{Q}\cong L$ is also diagonal.
If we denote by $M_{v}$ and $M^{\prime}_{v}$ the matrices representing $f$ and $f^{\prime}$ (as above), we conclude, as in the proof of Lemma 13, that the corresponding endormorphisms $\phi$ and $\phi^{\prime}$ on $H^{1}_{v}(\mathcal{E}_{2,s})$ and $H^{1}_{v}(\mathcal{E}_{3,s})$ are represented by $d_{1}N_{0}M_{v}$ and $d_{2}N^{\prime}_{0}M^{\prime}_{v}$, respectively. Hence, we conclude that $M_{v}$ and $M^{\prime}_{v}$ are diagonal.
Again, as at (3), we obtain
$\Lambda_{v}$ $:=(F_{2,s})^{-1}A^{T}F_{1,s}=\Omega_{v,0}M_{v}\Omega_{v,0}^{-1};$
$\Lambda^{\prime}_{v}$ $:=(F_{3,s})^{-1}(A^{\prime})^{T}F_{1,s}=\Omega_{v,0}M^{\prime}_{v}\Omega_{v,0}^{-1}.$
Since $M_{v}$ and $M^{\prime}_{v}$ are diagonal, $\Lambda_{v}$ and $\Lambda^{\prime}_{v}$ commute, so we obtain
(7) $\Lambda_{v,21}\Lambda^{\prime}_{v,12}=\Lambda_{v,12}\Lambda^{\prime}_{v,21}.$
Each entry of $\Lambda_{v}$ or $\Lambda^{\prime}_{v}$ is the evaluation at $(F_{imn}^{v\textrm{-}\mathrm{an}}(x(s)^{v\textrm{-}\mathrm{an}})_{1\leq i\leq 3,\,1\leq m,n\leq 2}$ of a polynomial of degree $4$, with coefficients in $K$, which is independent of $v$. Thus, (7) gives us the desired $v$-adic relation between evaluations of $\mathcal{G}$, independent of $v$.
Again, the fact that these relations are not in $I$ can be seen by elementary computation. ∎
4. Proving Theorems 5 and 6
In this section, we prove the main height bounds, namely, Theorems 5 and 6. Before doing so, we collect some general constructions that we will use in the proofs.
4.1. Elliptic curve schemes
Let $\mathcal{E}\to S:=\mathbb{A}^{1}_{\mathbb{Q}}\setminus\{0,1728\}$ denote the “$j$-family” of elliptic curves defined by the equation
$y^{2}+xy=x^{3}-\frac{36}{j-1728}x-\frac{1}{j-1728}.$
As per the name, for $j_{0}\in\mathbb{A}^{1}(\overline{\mathbb{Q}})\setminus\{0,1728\}$, the fiber $\mathcal{E}_{j_{0}}$ is an elliptic curve with $j$-invariant $j(\mathcal{E}_{j_{0}})=j_{0}$. The fiber at $\infty$ of the Zariski closure of $\mathcal{E}$ in $\mathbb{P}^{1}\times\mathbb{P}^{2}$ is a nodal cubic, so the fibre at $\infty$ of the connected Néron model of $\mathcal{E}$ over $S^{\prime}:=\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1728\}$ is isomorphic to $\mathbb{G}_{m}$.
Now let $n\in\mathbb{N}$ and suppose that $C\subset S_{K}^{n}$ is a geometrically irreducible algebraic curve defined over a number field $K$. For $i\in\{1,\ldots,n\}$, we obtain elliptic schemes $\mathcal{E}_{i}\to C$ by pulling back $\mathcal{E}\to S$ along the co-ordinate projections $C\to S_{K}$. We refer to the $\mathcal{E}_{i}\to C$ as the standard elliptic curve schemes on $C$.
4.2. Covering data
Consider the situation in Section 4.1 and let $C^{\prime}$ denote a smooth compactification of $C$. After possibly replacing $C^{\prime}$ with a finite étale cover, we may assume that the connected Néron model $\mathcal{E}^{\prime}_{i}$ of $\mathcal{E}_{i}$ over $C^{\prime}$ is semiabelian [Sil86, Prop. VII.5.4].
Let $s_{0}\in C^{\prime}(K)$. By [DO23, Lemma 6.6], after possibly replacing $K$ with a finite extension, we obtain
  • (a)
    a smooth projective geometrically irreducible algebraic curve $\tilde{C}^{\prime}$ over $K$;
  • (b)
    a non-constant morphism $\nu:\tilde{C}^{\prime}\to C^{\prime}$ over $K$;
  • (c)
[…]
  • (d)
    a regular $\mathcal{O}_{K}$-model $\tilde{\mathfrak{C}}^{\prime}$ of $\tilde{C}^{\prime}$;
  • (e)
    semiabelian schemes $\mathfrak{G}^{\prime}_{i}\to\tilde{\mathfrak{C}}^{\prime}$
such that
  • (i)
    every point $s\in\tilde{C}^{\prime}(\overline{K})$ satisfying $x(s)=0$ is a simple zero of $x$ and satisfies $\nu(s)=s_{0}$;
  • (ii)
    $x:\tilde{C}^{\prime}\to\mathbb{P}^{1}$ is Galois, which is to say, $\operatorname{Aut}_{x}(\tilde{C}^{\prime})=\{\sigma\in\operatorname{Aut}(\tilde{C}^{\prime}):x\circ\sigma=x\}$ acts transitively on each $\overline{K}$-fibre of $x$;
  • (iii)
    $\mathfrak{G}^{\prime}_{i,K}\cong\tilde{\mathcal{E}}^{\prime}_{i}:=\tilde{C}^{\prime}\times_{C^{\prime}}\mathcal{E}^{\prime}_{i}$.
[…]
By [DO23, Lemma 6.2], there exists another regular $\mathcal{O}_{K}$-model $\tilde{\mathfrak{C}}^{\prime\prime}$ of $\tilde{C}^{\prime}$ such that the $\sigma_{j}$ extend to morphisms $\tilde{\mathfrak{C}}^{\prime\prime}\to\tilde{\mathfrak{C}}$$^{\prime}$. Let $\tilde{C}=\nu^{-1}(C)$. We refer to
$\mathcal{D}:=(\mathfrak{G}^{\prime}_{1},\ldots,\mathfrak{G}^{\prime}_{n},\tilde{\mathfrak{C}}^{\prime\prime},\tilde{\mathfrak{C}}^{\prime},\tilde{C}^{\prime},\tilde{C},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
as a covering datum for $(C^{\prime},s_{0})$.
4.3. Pullback representatives
Continue from the situation obtained in Section 4.2. For $I\subset\{1,\ldots,n\}$ and $\lambda=(j,i)\in\{1,\dotsc,\ell\}\times I$, we write $\mathfrak{G}^{\prime}_{\lambda}:=\sigma^{*}_{j}\mathfrak{G}^{\prime}_{i}$ and we write $\mathcal{E}^{\prime}_{\lambda}:=\mathfrak{G}^{\prime}_{\lambda,K}$. Then $\mathfrak{G}^{\prime}_{\lambda}$ is a scheme over $\tilde{\mathfrak{C}}^{\prime\prime}$ and $\mathcal{E}^{\prime}_{\lambda}$ is a scheme over $\tilde{\mathfrak{C}}^{\prime\prime}_{K}=\tilde{C}^{\prime}$.
Let $\bar{\eta}$ denote a geometric generic point of $\tilde{C}^{\prime}$ and define an equivalence relation $\sim$ on $\{1,\dotsc,\ell\}\times I$ by the condition
[…]
Let $\Lambda$ denote a set of representatives for the induced equivalence classes. We refer to $\Lambda$ as a set of pullback representatives for $(\mathcal{D},I)$. We will write $[j,i]$ for the unique element of $\Lambda$ equivalent to $(j,i)$.
4.4. Neighbourhood systems
Continue from the situation obtained in Section 4.3. To simplify notation, relabel $C^{\prime}:=\tilde{C}^{\prime}$. For each archimedean (resp. non-archimedean) place $v$ of $K$, apply [DO22, Lemma 5.5] (resp. Proposition 21) to $C^{\prime}$ and $x\in K(C^{\prime})$, thereby obtaining real numbers $r_{v}>0$ and open subspaces $U_{v,1},\dotsc,U_{v,\ell}$ of $C^{\prime{v\textrm{-}\mathrm{an}}}$ with the properties (i)–(v) of [DO22, Lemma 5.5] (resp. Proposition 21). Let $\zeta\in K^{\times}$ such that $|\zeta|_{v}\leq r_{v}$ for the finitely many places $v$ of $K$ for which $r_{v}<1$ and let $H$ denote the G-function $\zeta/(\zeta-X)$. We refer to $((r_{v},U_{v,1},\dotsc,U_{v,\ell})_{v},H)$ as a neighbourhood system for $(C^{\prime},x)$.
4.5. Proof of Theorem 5
Consider the situation in Theorem 5 and let $K$ denote the field of definition of $C$.
4.5.1. Reductions
After possibly relabeling the coordinates, we can assume that $\overline{C}\setminus C$ contains a point $s_{0}:=(b_{1},b_{2},\infty)$ with $(b_{1},b_{2})\in Y(1)^{2}$ modular. After possibly removing finitely many points from $C$, namely, those belonging to the special hyperplanes defined by setting a coordinate to 0 or $1728$, we can assume that $C$ is contained in $S^{3}_{K}$. After possibly replacing $K$ with a finite extension and $\overline{C}$ with a finite étale cover, we obtain a covering datum
$\mathcal{D}:=(\mathfrak{G}^{\prime}_{1},\mathfrak{G}^{\prime}_{2},\mathfrak{G}^{\prime}_{3},\tilde{\mathfrak{C}}^{\prime\prime},\tilde{\mathfrak{C}}^{\prime},\tilde{C}^{\prime},\tilde{C},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
[…]
4.5.2. Setup
For $I:=\{1,2\}$, let $\Lambda\subset\{1,\ldots,\ell\}\times I$ be a set of pullback representatives for $(\mathcal{D},I)$. As in [DO22, Rem. 5.3], we have $j,1]\neq[j,2]$ for all $j\in\{1,\ldots,\ell\}$. Let $\tilde{\mathfrak{C}}\subset\tilde{\mathfrak{C}}$$^{\prime\prime}$ denote the maximal Zariski open subset such that $\mathfrak{G}_{\lambda}:=\mathfrak{G}^{\prime}_{\lambda}\times_{\tilde{\mathfrak{C}}^{\prime\prime}}\tilde{\mathfrak{C}}$ is an elliptic scheme for all $\lambda\in\Lambda$ and replace $\tilde{C}$ with $\tilde{\mathfrak{C}}_{K}$ (note that we still have $s_{j}\in\tilde{C}(K)$ for all $j\in\{1,\ldots,\ell\})$. For $\lambda\in\Lambda$, define $\mathcal{E}_{\lambda}:=\mathcal{E}^{\prime}_{\lambda}|_{\tilde{C}}$.
To simplify notation, remove all $\tilde{\cdot}$ decorations. By [DO22, Lemma 5.4], for any $j\in\{1,\ldots,\ell\}$, there exists an isogeny
$\mathcal{E}_{[j,1],s_{1}}=\mathcal{E}^{\prime}_{[j,1],s_{1}}\to(\sigma^{*}_{j}\mathcal{E}^{\prime}_{1})_{s_{1}}=\mathcal{E}^{\prime}_{1,s_{j}}\to\mathcal{E}^{\prime}_{2,s_{j}}=(\sigma^{*}_{j}\mathcal{E}^{\prime}_{2})_{s_{1}}\to\mathcal{E}^{\prime}_{[j,2],s_{1}}=\mathcal{E}_{[j,2],s_{1}}$
of elliptic curves, which we denote $f_{j}$. We choose the $f_{j}$ compatibly, so that, if $j,i]=[j^{\prime},i]$, then the induced diagram of isogenies commutes. Furthermore, after possibly replacing $K$ with a finite extension, we may assume that the $f_{j}$ are defined over $K$ and, after possibly removing finitely many points from $C$, that $H^{1}_{DR}(\mathcal{E}_{[j,i]}/C)$ is free. For each $\lambda=[j,i]\in\Lambda$, choose bases $\{\omega_{\lambda},\eta_{\lambda}\}$ for $H^{1}_{DR}(\mathcal{E}_{\lambda}/C)$ as in Section 3, with $s_{1}\in C(K)$ playing the role of $s_{0}$ and $f_{j}$ the role of $f_{0}$. (This is possible due to the compatibility of the $f_{j}$.)
Let $F_{\lambda mn}\in K[[X]]$ ($1\leq m,n\leq 2$) denote the G-functions associated with $\mathcal{E}_{\lambda}\to C$, $x$ and $\{\omega_{\lambda},\eta_{\lambda}\}$ (see Section 2.6). We define
$\mathcal{G}=\{F_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2\},$
and, for each place $v$ of $K$, we define
$R_{v}=\min\{1,R(F^{v\textrm{-}\mathrm{an}}):F\in\mathcal{G}\}.$
Let $((r_{v},U_{v,1},\dotsc,U_{v,\ell})_{v},H)$ denote a neighbourhood system for $(C^{\prime},x)$ and define $\mathcal{G}_{H}:=\mathcal{G}\cup\{H\}$. For each $v\in\Sigma_{K}$, define
$R^{\leq}_{v}:=\min\{R_{v},R(H^{v\textrm{-}\mathrm{an}})\}\leq r_{v},$
and, for each $j\in\{1,\ldots,\ell\}$, let $U^{\prime}_{v,j}=(x^{v\textrm{-}\mathrm{an}}|_{U_{v,j}})^{-1}(D(0,R^{\leq}_{v},K_{v}))$. By Lemma 9, after possibly modifying $H$ (in other words, altering finitely many of the $r_{v}$), we can ensure that, for any $j\in\{1,\ldots,\ell\}$, the data of
$C^{\prime},\mathfrak{C}^{\prime\prime},\mathfrak{G}^{\prime}_{\lambda},C,\mathcal{E}_{\lambda},\mathfrak{C},\mathfrak{G}_{\lambda},s_{1},f_{j},\{\omega_{\lambda},\eta_{\lambda}\},x,F_{\lambda mn},\Delta_{v}:=U^{\prime}_{v,1},$
for $\lambda=[j,1],[j,2]$ and $v\in\Sigma_{K}$, satisfy the assumptions of Section 3.
4.5.3. Constructing relations
Consider $s\in C(\overline{\mathbb{Q}})$ as above and replace $K$ with $K(s)$. For any $j\in\{1,\ldots,\ell\}$, we obtain an isogeny
$\mathcal{E}_{[j,1],\sigma_{j}^{-1}(s)}\to(\sigma^{*}_{j}\mathcal{E}_{1})_{\sigma^{-1}_{j}(s)}=\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}=(\sigma^{*}_{j}\mathcal{E}_{2})_{\sigma^{-1}_{j}(s)}\to\mathcal{E}_{[j,2],\sigma_{j}^{-1}(s)},$
the outer isogenies being afforded to us, again, by [DO22, Lemma 5.4]. Moreover, for any $v\in\Sigma_{K}$ such that $|x(s)|_{v}<R^{\leq}_{v}$, we have $s^{v\textrm{-}\mathrm{an}}\in U^{\prime}_{v,j}$ for some $j\in\{1,\ldots,\ell\}$. As in [DO22, Rem. 5.6], we deduce that $\sigma^{-1}_{j}(s)^{{v\textrm{-}\mathrm{an}}}\in U^{\prime}_{v,1}=\Delta_{v}$. Hence, by Proposition 11, there exists a $v$-adic relation $Q_{v}$ of degree at most $2$ between the evaluations at $x(\sigma^{-1}_{j}(s))=x(s)$ of the $F_{[i,j]mn}$ for $i,m,n\in\{1,2\}$ which is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i,m,n\leq 2]$ generated by the elements
$\det((X_{1mn})_{mn})-\det((X_{2mn})_{mn}).$
Let $\Sigma_{s}$ denote the set of $v\in\Sigma_{K}$ satisfying $|x(s)|_{v}<R^{\leq}_{v}$ and let
$Q:=\prod_{v\in\Sigma_{s}}Q_{v}.$
Then $Q$ is a global relation between the evaluations at $x(s)$ of the elements of $\mathcal{G}$ (and, hence, $\mathcal{G}_{H}$). To constrain its degree, we observe that $v\in\Sigma_{s}$ implies that $v\in\Sigma_{K,\infty}$ or $v\in\Sigma_{K,f}$ and, writing $\mathfrak{s}_{1}$, $\mathfrak{s}_{1}(j)$, and $\mathfrak{s}(j)$ for the canonical extensions of $s_{1}$, $\sigma^{-1}_{j}(s_{1})$, and $\sigma^{-1}_{j}(s_{1})$, respectively, we have an isogeny
$\mathfrak{G}^{\prime}_{1,\mathfrak{s}_{1},k_{v}}=\mathfrak{G}^{\prime}_{[j,1],\mathfrak{s}_{1}(j),k_{v}}\cong\mathfrak{G}^{\prime}_{[j,1],\mathfrak{s}(j),k_{v}}\to\mathfrak{G}^{\prime}_{[j,3],\mathfrak{s}(j),k_{v}}\cong\mathfrak{G}^{\prime}_{[j,3],\mathfrak{s}_{1}(j),k_{v}}=\mathfrak{G}^{\prime}_{3,\mathfrak{s}_{1},k_{v}}=\mathbb{G}_{m,k_{v}},$
where we use the fact that $\sigma^{-1}_{j}(s)^{{v\textrm{-}\mathrm{an}}}\in\Delta_{v}$. In other words, $v\in\Sigma_{s}$ implies that $v\in\Sigma_{K,\infty}$ or $v\in\Sigma_{K,f}$ and $v$ is a place of bad reduction for $\mathcal{E}_{1,s_{0}}$.
[…]
4.5.4. Nontriviality
By [DO22, Lemma 2.9], it suffices to show that $Q$ is a non-trivial relation between the elements of $\mathcal{G}$ (as opposed to $\mathcal{G}_{H}$). Therefore, let $J$ denote the (homogeneous) ideal of
$\overline{\mathbb{Q}}[X][X_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2\}]$
comprising the functional relations between the elements of $\mathcal{G}$. We need to show that $Q\notin J$. By the following lemma, it suffices to show that $Q_{v}\notin J$ for all $v\in\Sigma_{s}$.
Lemma 15.
The ideal $J$ is prime.
Proof.
By definition, $J$ is the homogeneous part of the kernel of the homomorphism
[…]
defined by $X_{\lambda mn}\mapsto F_{\lambda mn}$. Since the latter is clearly prime, so too is $J$ by [Sta, 00JM, Lemma 10.57.7]. ∎
To ease notation, we set $k:=|\Lambda|$. Choose $v\in\Sigma_{K,\infty}$. The $F_{\lambda mn}$ give rise to a function $D(0,R^{\leq}_{v},\mathbb{C})\to\mathbf{SL}^{k}_{2}(\mathbb{C})$, and we let $\Gamma$ denote the graph of this function.
Lemma 16.
The set $\Gamma$ is $\mathbb{C}$-Zariski dense in $\mathbb{A}^{1}\times\mathbf{SL}_{2}^{k}$.
Proof.
It suffices to show that
$\operatorname{trdeg}_{\mathbb{C}(X)}\mathbb{C}(X)(F_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2)\geq 3k.$
To that end, denote by $\Omega_{v}:\Delta_{v}\to\mathrm{M}_{2g}(\mathbb{C})$ the function giving rise to the $F_{\lambda mn}$, as in Section 2.4, and we write $\Omega_{v\lambda mn}$ for its components. Then the previous inequality is clearly implied by
$\operatorname{trdeg}_{\mathbb{C}(C)}\mathbb{C}(C)(\Omega_{v\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2)\geq 3k.$
As in Remark 7, if we define $\Omega_{v}$ with suitable horizontal bases for $H^{1}_{DR}(\mathcal{E}_{\lambda}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$, the $\Omega_{v\lambda mn}$ are the (non-zero) entries of a period matrix for $\mathcal{E}_{\lambda}\to C$. Therefore, the geometric André–Grothendieck period conjecture, proved in this case by Ayoub [Ayo15] and Nori (unpublished)—see [BT25, Thm. 1.1] and the remarks thereafter—yields exactly this inequality. ∎
By definition, we have $\Gamma\subset V(J)$. Thus, since $J$ is homogeneous, Lemma 16 implies that
$V(J)=\mathbb{A}^{1}\times\{(g_{1},\ldots,g_{k})\in\mathbf{GL}_{2}^{k}(\mathbb{C}):\det(g_{1})=\ldots=\det(g_{k})\}.$
Since $J$ is prime and therefore radical, we deduce that $J$ is the ideal generated by the elements
$\det((X_{\lambda_{1}mn})_{mn})-\det((X_{\lambda_{2}mn})_{mn})$
for $\lambda_{1},\lambda_{2}\in\Lambda$.
Now let $v\in\Sigma_{s}$. The fact that $Q_{v}\notin J$ now follows from the fact that $Q_{v}$ is not contained in the ideal $I$ of Proposition 11 (by faithful flatness of a polynomial ring over its ring of coefficients, for example). ∎
4.6. Proof of Theorem 6
Consider the situation in Theorem 6. In particular, let $s_{0}:=(b_{1},b_{2},b_{3})\in C$ be the modular point referred to in the statement, and let $K$ denote a number field over which $C$ and $s_{0}$ are defined. Let $\epsilon>0$.
4.6.1. Reductions
After removing at most finitely many points from $C$, we can assume that $C$ is contained in $S^{3}_{K}$ and we denote by $\mathcal{E}_{i}\to C$ the standard elliptic curve schemes on $C$. After possibly replacing $K$ with a finite extension and $\overline{C}$ with a finite étale cover, we obtain a covering datum
$\mathcal{D}:=(\mathfrak{G}^{\prime}_{1},\mathfrak{G}^{\prime}_{2},\mathfrak{G}^{\prime}_{3},\tilde{\mathfrak{C}}^{\prime\prime},\tilde{\mathfrak{C}}^{\prime},\tilde{C}^{\prime},\tilde{C},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
for $(\overline{C},s_{0})$, and we fix a Weil height $h$ on $\tilde{C}^{\prime}$. As before, it suffices for us to show that there exist constants $c_{9}$, $c_{10}$} and $\delta$ such that, for any modular point $s:=(x_{1},x_{2},x_{3})\in C$ of supersingular exponent $\delta$, and any $\tilde{s}\in\nu^{-1}(s)\in\tilde{C}$, we have
[…]
4.6.2. Setup
We copy the notations and repeat the constructions of Section 4.5, this time with $I:=\{1,2,3\}$. In particular, for any $s=(x_{1},x_{2},x_{3})$ and $\tilde{s}$, as above, and any $w\in\Sigma_{K(\tilde{s}),f}$ lying above any $v\in\Sigma_{K,f}$, the condition $|x(\tilde{s})|_{w}<R^{\leq}_{v}$ implies
$\mathfrak{G}^{\prime}_{[j,i],\mathfrak{s}(j),k_{w}}\cong\mathfrak{G}^{\prime}_{[j,i],\mathfrak{s}_{1}(j),k_{w}}$
for all $j\in\{1,\ldots,\ell\}$ and $i\in I$. It follows that, for any $i\in I$, the $j$-invariants $x_{i}$ and $b_{i}$ are congruent modulo the maximal ideal of $\mathcal{O}_{K(\tilde{s}),w}$, which is to say, $|x_{i}-b_{i}|_{w}<1$.
4.6.3. Constructing relations
[…]
$Q^{\rm nord}:=\prod_{v\in\Sigma^{\rm nord}_{\tilde{s}}}Q_{v}$
with $Q_{v}$ a relation of degree at most $2$ and $\Sigma^{\rm nord}_{\tilde{s}}\subset\Sigma_{\tilde{s}}$ the complement of the set of finite places $v\in\Sigma_{\tilde{s}}$ for which the $\mathcal{E}_{i,s}$ have good ordinary reduction. The fact that $Q$ is nontrivial is proved exactly as in Section 4.5.4, hence, we omit it.
[…]
$|\Sigma^{\rm sup}_{\tilde{s}}|\leq\Delta(s)^{\delta}$
and so the result follows from [And89, Thm. 5.2] provided $3(6\ell-1)\delta<1/2$. Rearranging, this is equivalent to
$\delta<6^{-1}(6\ell-1)^{-1}$
and, since $\ell$ depends only on $C$ and $s_{0}$, the proof is complete. ∎
5. Proving Theorems 3 and 4
The proofs of Theorems 3 and 4 follow the Pila–Zannier strategy. In fact, following word-for-word the proof of [HP12, Prop. 5.1], it suffices to obtain the following large Galois orbit results, which serve as substitutes for [HP12, Lemma 4.2].
Theorem 17.
Let $C$ satisfy the conditions of Theorem 3. There exist constants $c_{12}$ and $c_{13}$ such that, for any point $(x_{1},x_{2},x_{3})$ as in the statement of Conjecture 2,
[…]
Theorem 18.
Let $C$ and $(b_{1},b_{2},b_{3})$ be as in Theorem 4. There exist constants $c_{14}$ and $c_{15}$ such that, for any point $(x_{1},x_{2},x_{3})$ as in the statement of Theorem 4,
[…]
The proofs of these theorems follow almost exactly the proof of [DO22, Prop. 5.15], which is essentially an extract from the proof of [HP12, Lemma 4.2]. The key input is a Masser–Wüstholz-type isogeny estimate—we cite the (effective) bound of Gaudron–Rémond [GR14, Thm. 1.4].
Appendix: rigid analytic neighbourhoods
We establish a version of [DO22, Lemma 5.5], written in terms of rigid geometry instead of only $\mathbb{C}_{p}$-points, and thereby give a rigid analytic proof of [DO22, Lemma 5.5]. We also establish a version which talks about the neighbourhood of a single point where $x$ has a simple zero, instead of requiring all zeroes of $x$ to be simple.
First we check that an étale morphism of rigid spaces restricts to an isomorphism on some open neighbourhood of each $K$-point. (Note that it is not true in general that étale morphisms are local isomorphisms – the open neighbourhoods on which they are isomorphisms need not form an admissible covering.)
Lemma 19.
Let $K$ be a non-archimedean field of characteristic zero. Let $X$ be a quasi-separated rigid space over $K$. Let $f\colon X\to\bar{D}(0,r,K)$ be a finite étale morphism from $X$ to the closed disc of radius $r$. Suppose there is a $K$-point $x_{0}\in X(K)$ such that $f(x_{0})=0$. Then there exists an open subspace $Z\subset X$ containing $x_{0}$ and a positive real number $r^{\prime}\leq r$ such that $f|_{Z}\colon Z\to\bar{D}(0,r^{\prime},K)$ is an isomorphism of rigid spaces.
Proof.
Since $f$ is étale, it is flat, so its image is open in $D(0,r,K)$. In particular, the image of $f$ contains some disc $D(0,R,K)$. So, after replacing $X$ by $f^{-1}(D(0,R,K))$ and $r$ by $R$, we may assume that $f$ is surjective (i.e. it is a finite étale covering).
Let $L$ denote the completion of the algebraic closure of $K$. By [Lüt93, Thm. 2.1], after extending scalars to $L$, the finite étale covering $f_{L}\colon X_{L}\to\bar{D}(0,r,L)$ splits over $\bar{D}(0,r^{\prime},L)$ for some $r^{\prime}$ with $0<r^{\prime}\leq r$. In other words, letting $Y=f^{-1}(\bar{D}(0,r,k))$, we have that $Y_{L}$ is the disjoint union of subspaces $Y_{1},\dotsc,Y_{\ell}$ such that $f_{L}|_{Y_{i}}\colon Y_{i}\to\bar{D}(0,r^{\prime},L)$ is an isomorphism for each $i$.
Let $Z$ denote the connected component of $Y$ which contains $x_{0}$. Since $Z$ is connected and it has a $K$-point $x_{0}$, $Z_{L}$ is connected [Con99, Thm. 3.2.1]. Hence $Z_{L}$ is a connected component of $Y_{L}$, that is, $Z_{L}=Y_{i}$ for some $i$. Thus $f_{L}|_{Z_{L}}\colon Z_{L}\to\bar{D}(0,r^{\prime},L)$ is an isomorphism, so $f|_{Z}\colon Z\to\bar{D}(0,r^{\prime},k)$ is an isomorphism [Con06, Thm. A.2.4]. ∎
Proposition 20.
Let $C$ be a smooth irreducible algebraic curve over a number field $K$. Let $x\in K(C)$ be a rational function. Let $s_{0}\in C(K)$ be a simple zero of $x$.
For each $v\in\Sigma_{K,f}$, there exists a real number $r_{v}>0$ and a rigid open subspace $U_{v}\subset C^{v\textrm{-}\mathrm{an}}$ with the following properties:
  1. (i)
    $r_{v}\geq 1$ for almost all $v\in\Sigma_{K,f}$;
  2. (ii)
    $s_{0}^{v\textrm{-}\mathrm{an}}\in U_{v}$;
  3. (iii)
    for each $v\in\Sigma_{K,f}$, the morphism $x^{v\textrm{-}\mathrm{an}}$ restricts to an isomorphism of rigid spaces from $U_{v}$ to the rigid open disc $D(0,r_{v},K_{v})$;
  4. (iv)
    for every rational function $f\in K(C)$ which is regular at $s_{0}$, if $\hat{f}\in K[\![X]\!]$ denotes the Taylor series of $f$ around $s_{0}$ in terms of the local parameter $x$, then, for all $s\in U_{v}$ satisfying $\lvert x^{v\textrm{-}\mathrm{an}}(s)\rvert<R(\hat{f}^{v\textrm{-}\mathrm{an}})$, we have $\hat{f}^{v\textrm{-}\mathrm{an}}(x^{v\textrm{-}\mathrm{an}}(s))=f^{v\textrm{-}\mathrm{an}}(s)$.
Proof.
By standard facts about algebraic curves, $C$ is isomorphic to a Zariski open subset of some smooth projective curve over $K$. The conclusions of the proposition are local around a given point of $C$, so we may replace $C$ by this smooth projective curve.
Consider $x$ as a morphism $C\to\mathbb{P}^{1}_{K}$. By [DO23, Lemma 6.1], there exists a regular $\mathcal{O}_{K}$-model $\mathfrak{C}$ of $C$ such that $x$ extends to a morphism of $\mathcal{O}_{K}$-schemes $\xi\colon\mathfrak{C}\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$. Since $x$ is a non-constant morphism between integral curves, it is flat. Since $\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$ is also flat, the fibre-wise criterion of flatness establishes that $\xi\colon\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$ is flat at all points of the generic fibre.
By hypothesis, $x$ is unramified, hence étale, at $s_{0}$. Therefore, by [Gro67, Rmk. 17.8.3], $\xi$ is étale at $s_{0}$. It follows that $\xi$ is étale on a non-empty Zariski open neighbourhood of $s_{0}$ in $\mathfrak{C}$.
By [DO23, Lemma 6.3], there exists a section $\mathfrak{s}_{0}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathfrak{C}$ which extends $s_{0}\colon\operatorname{Spec}(K)\to C$. It follows that there exists a dense open subset $\mathfrak{V}\subset\operatorname{Spec}(\mathcal{O}_{K})$ such that $\xi$ is étale on $\mathfrak{s}_{0}(\mathfrak{V})$.
We shall give two constructions of open subspaces $U_{v}\subset C^{v\textrm{-}\mathrm{an}}$ satisfying (ii) and (iii), depending on whether $\mathfrak{p}_{v}$ lies in $|\mathfrak{V}|$ or not. The construction for primes in $|\mathfrak{V}|$ gives $r_{v}=1$, while the construction for other primes may give $r_{v}<1$. This is sufficient to ensure that (i) holds.
First consider $v\in\Sigma_{K,f}$ such that $\mathfrak{p}:=\mathfrak{p}_{v}\in|\mathfrak{V}|$. Let $\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}$ denote the formal completion of $\mathfrak{C}$ along $\mathfrak{s}_{0}(\mathfrak{p})$ and let $\mathbb{P}^{1}_{\mathfrak{p},\mathrm{for}{\dagger}}$ denote the formal completion of $\mathbb{P}^{1}_{\mathcal{O}_{K}}$ along $\mathfrak{z}_{0}(\mathfrak{p})$, where $\mathfrak{z}_{0}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$ denotes the zero section. Since $\xi$ is étale at $\mathfrak{s}_{0}(\mathfrak{p})$, it induces an isomorphism of formal schemes $\xi_{\mathfrak{p},\mathrm{for}{\dagger}}\colon\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}\to\mathbb{P}^{1}_{\mathfrak{p},\mathrm{for}{\dagger}}$. Applying Berthelot’s rigid generic fibre functor, we obtain an isomorphism of rigid spaces $\xi_{\mathfrak{p}}^{\mathrm{rig}}\colon\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}^{\mathrm{rig}}\to(\mathbb{P}^{1}_{\mathfrak{p},\mathrm{for}{\dagger}})^{\mathrm{rig}}$. Here, $\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}^{\mathrm{rig}}$ is a rigid open subspace of $C^{v\textrm{-}\mathrm{an}}$, and $(\mathbb{P}^{1}_{\mathfrak{p},\mathrm{for}{\dagger}})^{\mathrm{rig}}=\operatorname{Spf}(\mathcal{O}_{K,\mathfrak{p}}[\![X]\!],(\mathfrak{p},X))^{\mathrm{rig}}$ is the rigid open unit disc. Thus, if we define $U_{v}$ to be $\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}^{\mathrm{rig}}$, then $U_{v}$ satisfies (ii) and (iii) with $r_{v}=1$.
Now consider $v\in\Sigma_{K,f}$ such that $\mathfrak{p}_{v}\notin|\mathfrak{V}|$ (this construction applies for all $v\in\Sigma_{K,f}$, but we use the previous construction for $v$ such that $\mathfrak{p}_{v}\in|\mathfrak{V}|$). Since $x$ is finite and étale on a Zariski open neighbourhood of $s_{0}$ in $C$, there is a rigid open neighbourhood $V_{v}$ of $x$ in $C^{v\textrm{-}\mathrm{an}}$ such that $x^{v\textrm{-}\mathrm{an}}|_{V_{v}}$ is finite étale. After shrinking $V_{v}$, we may assume that the image of $V_{v}$ is contained in a disc around 0. By Lemma 19, there exists an open subspace $U_{v}\subset V_{v}$ containing $s_{0}^{v\textrm{-}\mathrm{an}}$ such that $x^{v\textrm{-}\mathrm{an}}|_{U_{v}}$ is an isomorphism of rigid spaces from $U_{v}$ to a closed rigid disc $\bar{D}(0,r_{v},K_{v})$. After shrinking $U_{v}$, we may instead arrange that $x^{v\textrm{-}\mathrm{an}}|_{U_{v}}$ is an isomorphism onto the open rigid disc $D(0,r_{v},K_{v})$. Thus $U_{v}$ satisfies (ii) and (iii).
Finally we verify (iv) for any $U_{v}$ satisfying (ii) and (iii). Since the analytification functor induces a morphism of locally G-ringed spaces $(C^{v\textrm{-}\mathrm{an}},\mathcal{O}_{C^{v\textrm{-}\mathrm{an}}})\to(C_{K_{v}},\mathcal{O}_{C_{K_{v}}})$, the (algebraic) Taylor series of $f$ with respect to $x$ is the same as the (rigid analytic) Taylor series of $f^{v\textrm{-}\mathrm{an}}$ with respect to $x^{v\textrm{-}\mathrm{an}}$. Since the Taylor series of a rigid function $g$ on a disc converges to $g$ inside its radius of convergence, (iv) holds. ∎
Now we prove a rigid analogue of [DO22, Lemma 5.5].
Proposition 21.
Let $C$ be a smooth algebraic curve over a number field $K$. Let $x\in K(C)$ be a rational function of degree $\ell$, such that $x$ has $\ell$ distinct unramified zeros $s_{1},\dotsc,s_{\ell}\in C(K)$.
For each $v\in\Sigma_{K,f}$, there exists a real number $r_{v}>0$ and rigid open subspaces $U_{v,1},\dotsc,U_{v,\ell}\subset C^{v\textrm{-}\mathrm{an}}$ with the following properties:
  1. (i)
    $r_{v}\geq 1$ for almost all $v\in\Sigma_{K,f}$;
  2. (ii)
    $s_{k}^{v\textrm{-}\mathrm{an}}\in U_{v,k}$;
  3. (iii)
    for each $v\in\Sigma_{K,f}$, the spaces $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint, and form an admissible covering of the preimage under $x^{v\textrm{-}\mathrm{an}}$ of the rigid open disc $D(0,r_{v},K_{v})$;
  4. (iv)
    for each $v\in\Sigma_{K,f}$ and $k=1,\ldots,\ell$, the morphism $x^{v\textrm{-}\mathrm{an}}$ restricts to an isomorphism of rigid spaces from $U_{v,k}$ to the rigid open disc $D(0,r_{v},K_{v})$;
  5. (v)
    for every rational function $f\in K(C)$ which is regular at $s_{k}$, if $\hat{f}\in K[\![X]\!]$ denotes the Taylor series of $f$ around $s_{k}$ in terms of the local parameter $x$, then, for all $s\in U_{v,k}$ satisfying $\lvert x^{v\textrm{-}\mathrm{an}}(s)\rvert<R(\hat{f}^{v\textrm{-}\mathrm{an}})$, we have $\hat{f}^{v\textrm{-}\mathrm{an}}(x^{v\textrm{-}\mathrm{an}}(s))=f^{v\textrm{-}\mathrm{an}}(s)$.
Proof.
Apply Proposition 20 to each of $s_{1},\dotsc,s_{k}$, thereby yielding $r_{v,1},\dotsc,r_{v,\ell}$ and $U_{v,1},\dotsc,U_{v,\ell}$ for all $v\in\Sigma_{K,f}$. These satisfy (ii), (iv) and (v). However, for a given $v$, the subspaces $U_{v,1},\dotsc,U_{v,\ell}$ might not be pairwise disjoint.
As in the proof of Proposition 20, we may choose a regular model $\mathfrak{C}$ of $C$ such that $x$ extends to a morphism $\xi\colon\mathfrak{C}\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$, as well as sections $\mathfrak{s}_{1},\dotsc,\mathfrak{s}_{k}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathfrak{C}$ which extend $s_{1},\dotsc,s_{\ell}$. There is a dense open subset $\mathfrak{V}\subset\operatorname{Spec}(\mathcal{O}_{K})$ such that $\xi$ is étale on $\mathfrak{s}_{k}(\mathfrak{V})$ for all $k=1,\dotsc,\ell$. Since $s_{1},\dotsc,s_{k}$ are pairwise distinct, after shrinking $\mathfrak{V}$, we may also assume that $\mathfrak{s}_{j}(\mathfrak{p})\neq\mathfrak{s}_{k}(\mathfrak{p})$ for all $\mathfrak{p}\in|\mathfrak{V}|$ and all $j\neq k$.
We will show that, whenever $\mathfrak{p}_{v}\in|\mathfrak{V}|$, the $U_{v,1},\dotsc,U_{v,\ell}$ will be pairwise disjoint and form an admissible covering of their union. Indeed, let $\mathfrak{C}_{\mathfrak{p}}$ denote the fibre above $\mathfrak{p}:=\mathfrak{p}_{v}$ of $\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$. Let $\mathfrak{C}_{\mathfrak{p},\mathrm{for}}$ and $\mathfrak{C}_{\mathfrak{p},\mathrm{for},k{\dagger}}$ denote the formal completions of $\mathfrak{C}$ along $\mathfrak{C}_{\mathfrak{p}}$ and $\mathfrak{s}_{k}(\mathfrak{p})$ respectively. Since $\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$ is proper, $\mathfrak{C}_{\mathfrak{p}}^{\mathrm{rig}}\to C^{v\textrm{-}\mathrm{an}}$ is an isomorphism, so we obtain a reduction map $\operatorname{red}\colon C^{v\textrm{-}\mathrm{an}}\to\mathfrak{C}_{\mathfrak{p}}$. By construction, we have
$U_{v,k}=\mathfrak{C}_{\mathfrak{p},\mathrm{for},k{\dagger}}^{\mathrm{rig}}=\operatorname{red}^{-1}(\mathfrak{s}_{k}(\mathfrak{p})).$
Since $\mathfrak{s}_{1}(\mathfrak{p}),\dotsc,\mathfrak{s}_{\ell}(\mathfrak{p})$ are distinct (by our choice of $\mathfrak{p}\in|\mathfrak{V}|$), we deduce that $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint.
Consider the sets
$Z_{\mathfrak{p},k}:=\mathfrak{C}_{\mathfrak{p}}\setminus\{s_{j}:1\leq j\leq\ell,\,j\neq k\}$
for $1\leq k\leq\ell$, which form a Zariski open covering of $\mathfrak{C}_{\mathfrak{p}}$. Then the rigid spaces $\operatorname{red}^{-1}(Z_{\mathfrak{p},1}),\dotsc,\operatorname{red}^{-1}(Z_{\mathfrak{p},\ell})$ form an admissible covering of $C^{v\textrm{-}\mathrm{an}}$. Since $U_{v,k}\subset\operatorname{red}^{-1}(Z_{\mathfrak{p},k})$ and $U_{v,k}$ is disjoint from $\operatorname{red}^{-1}(Z_{\mathfrak{p},j})$ for $j\neq k$, it follows that $U_{v,1},\dotsc,U_{v,\ell}$ form an admissible covering of their union.
Meanwhile, if $\mathfrak{p}_{v}\not\in|\mathfrak{V}|$, we shrink $U_{v,1},\dotsc,U_{v,\ell}$ to make them disjoint and ensure that they form an admissible covering. To that end, choose $r^{\prime}_{v}<\min\{r_{v,1},\dotsc,r_{v,\ell}\}$ and let $U_{v,k}^{\prime}$ denote the preimage of the closed disc $\bar{D}(0,r_{v}^{\prime},K_{v})$ inside $U_{v,k}$. Then the rigid spaces $U_{v,k}^{\prime}$ are quasi-compact, so $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$ are quasi-compact for each $j,k$. Hence, if $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}\neq\emptyset$, the absolute value of the function $x^{v\textrm{-}\mathrm{an}}$ has a minimum value $R_{v,jk}$ on $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$. If $j\neq k$, then $s_{j}^{v\textrm{-}\mathrm{an}},s_{k}^{v\textrm{-}\mathrm{an}}\not\in U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$, so $R_{v,jk}>0$. Let
$r_{v}=\min\{r_{v}^{\prime},\,R_{v,jk}:j\neq k,U_{v,j}^{\prime}\cap U_{v,k}^{\prime}\neq\emptyset\}>0.$
By construction, after replacing $U_{v,k}$ by the preimage of $D(0,r_{v},K_{v})$ in itself, the sets $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint.
Since the sets $U_{v,k}^{\prime}$ are affinoid and there are finitely many of them, they form an admissible covering of their union. After the final replacement of $U_{v,k}$, we have $U_{v,k}\subset U_{v,k}^{\prime}$ and $U_{v,k}$ is disjoint from $U_{v,j}^{\prime}$ for $j\neq k$. Hence $U_{v,1},\dotsc,U_{v,\ell}$ form an admissible covering of their union.
Thus, for each $v\in\Sigma_{K,f}$, we have constructed $U_{v,1},\dotsc,U_{v,\ell}$ satisfying (ii), (iv) and (v), which are pairwise disjoint and form an admissible covering of their union. Furthermore, for all $\mathfrak{p}_{v}\in|\mathfrak{V}|$, we have used the sets $U_{v,k}$ from Lemma 19 without shrinking them, and these satisfy $r_{v}=1$. Thus (i) is satisfied.
To conclude, we note that every point of $(\mathbb{A}^{1})^{v\textrm{-}\mathrm{an}}$ has at most $\ell$ preimages under $x^{v\textrm{-}\mathrm{an}}$. On the other hand, since $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint, every point of $D(0,r_{v},K_{v})$ has $\ell$ distinct preimages in $U_{v,1}\cup\dotsb\cup U_{v,\ell}$. Hence the union $U_{v,1}\cup\dotsb\cup U_{v,\ell}$ is equal to $(x^{v\textrm{-}\mathrm{an}})^{-1}(D(0,r_{v},K_{v}))$. ∎